OpenStudy (anonymous):

solve for n in terms of c cn^20 = n^n

OpenStudy (mathstudent55):

You need to isolate c, so divide both sides by n^20.

OpenStudy (anonymous):

once I do that how can I get the n down from the exponent?

OpenStudy (anonymous):

c=n^(n-20)

OpenStudy (mathstudent55):

a^m/a^n = a^(m - n)

OpenStudy (anonymous):

I need to find the value of n in terms of c rather than the other way around

OpenStudy (mathstudent55):

You are correct.